In a grocery store there are 10 brands of canned tuna, 3 of which are on sale, a market goer takes 5 canned tuna at random, what is the probability that less than 3 are on sale?


1
Expert's answer
2021-05-17T17:14:01-0400

There are 10 cannned, 7 normal and 3 are from sale

The ways of selecting 5 canned out of total 10 canned is 10C5=252

The ways of selecting 1 from sale and 4 from normal=3C1*7C4=

The ways of selecting 2 from sale and 3 from normal=3C2*7C3=

The ways of selecting 0 from sales and 5 from normal=3C0*7C5

Required probability=(3C1*7C4+3C2*7C3+3C0*7C5)/10C5

(105+105+21)/252=0.9167